# WildAvatar: Web-scale In-the-wild Video Dataset for 3D Avatar Creation

### v1

@@ Abstract
Existing human datasets for avatar creation are typically limited to laboratory environments, wherein high-quality annotations (e.g., SMPL estimation from 3D scans or multi-view images) can be ideally provided. However, their annotating requirements are impractical for real-world images or videos, posing challenges toward real-world applications on current avatar creation methods. To this end, we propose the WildAvatar dataset, a web-scale in-the-wild human avatar creation dataset extracted from YouTube, with $10,000+$ different human subjects and scenes. WildAvatar is at least 10×10$\times$10 × richer than previous datasets for 3D human avatar creation. We evaluate several state-of-the-art avatar creation methods on our dataset, highlighting the unexplored challenges in real-world applications on avatar creation. We also demonstrate the potential for generalizability of avatar creation methods, when provided with data at scale. We will publicly release our data source links and annotations, to push forward 3D human avatar creation and other related fields for real-world applications.
@@ 3.2 Data Processing
Stage IV: Refining SMPL and Camera In-the-loop. The coarse SMPL estimated in Stage I may ignore the temporal consistency of human motions. Inspired by previous SMPL annotation pipelines [36, 2, 31], we further refine SMPL and camera parameters by smoothing these parameters across the whole video clip sequence via gradient descent. We also incorporate the estimated 2D keypoints and SAM masks into a refinement loop, providing additional supervision for precise SMPL and camera parameter estimation. Qualitative comparisons between coarse and refined SMPL and camera parameters can be found in the supplementary material.
@@ 4.2 Track A: Per-Subject Human Avatar Creation
Figure 4: Qualitative comparisons of popular per-subject methods on WildAvatar. NHP and NB, AN, AS, HN, IN, and GH denote Neural Human Rendering, NeuralBody, Animatable NeRF, Animatable SDF, HumanNeRF, InstantNVR, and GauHuman, respectively.
@@ 4.3 Track B: Generalizable Human Avatar Creation
Figure 5: Qualitative comparisons on the state-of-the-art generalizable avatar creation method [24]. “w/ WA” or “w/o WA” denotes training with or without WildAvatar, respectively.

### v2

@@ Abstract
Existing human datasets for avatar creation are typically limited to laboratory environments, wherein high-quality annotations (e.g., SMPL estimation from 3D scans or multi-view images) can be ideally provided. However, their annotating requirements are impractical for real-world images or videos, posing challenges toward real-world applications on current avatar creation methods. To this end, we propose the WildAvatar dataset, a web-scale in-the-wild human avatar creation dataset extracted from YouTube, with $10,000+$ different human subjects and scenes. WildAvatar is at least 10×10$\times$10 × richer than previous datasets for 3D human avatar creation. We evaluate several state-of-the-art avatar creation methods on our dataset, highlighting the unexplored challenges in real-world applications on avatar creation. We also demonstrate the potential for generalizability of avatar creation methods, when provided with data at scale. We publicly release our data source links and annotations, to push forward 3D human avatar creation and other related fields for real-world applications.
@@ 3.2 Data Processing
Stage IV: Refining SMPL and Camera In-the-loop. The coarse SMPL estimated in Stage I may ignore the temporal consistency of human motions. Inspired by previous SMPL annotation pipelines [37, 2, 32], we further refine SMPL and camera parameters by smoothing these parameters across the whole video clip sequence via gradient descent. We also incorporate the estimated 2D keypoints and SAM masks into a refinement loop, providing additional supervision for precise SMPL and camera parameter estimation. Qualitative comparisons between coarse and refined SMPL and camera parameters can be found in the Fig. A1.
@@ 4.2 Track A: Per-Subject Human Avatar Creation
Figure 4: Qualitative comparisons of popular per-subject methods on WildAvatar. NHP and NB, AN, AS, HN, IN, and GH denote Neural Human Rendering, NeuralBody, Animatable NeRF, Animatable SDF, HumanNeRF, InstantNVR, and GauHuman, respectively. Faces are blurred to protect privacy.
@@ 4.3 Track B: Generalizable Human Avatar Creation
Figure 5: Qualitative comparisons on the state-of-the-art generalizable avatar creation method [25]. “w/ WA” or “w/o WA” denotes training with or without WildAvatar, respectively. Faces are blurred to protect privacy.
@@ 5 Conclusions
Pipeline Efficiency. The efficiency of data collection and annotation is also crucial to facilitate data scale-up and application. Despite the complex multiple-stage design, our data-collecting pipeline only takes less than $200$ seconds to generate full annotations for an arbitrary in-the-wild video.
@@
Appendix A Details in Data Pipeline
We provide more details on the data processing pipeline in addition to the main paper.
Stage I: Human Bounding Box Detection and Tracking. We first obtain the bounding box of human subjects with an off-the-shelf state-of-the-art Yolo detection method [55]  [^2] The configuration of the detection model can be found here.. We only keep the video clip with at least one “person” instance with its detection threshold over $0.8$. The tracking step is based on finding the largest IOU overlay of bounding boxes among frames. We discard low-resolution human subjects whose bounding box areas are lower than $64\times 64$. To ensure the richness of the dataset, we only keep one “key subject” for each video, as clips of the same video may probably share the same key subject.
Stage II: Human Segmentation Mask Extraction. To prepare key point exemplars for the Segment Anything, we first obtain the 2D keypoints $J_{2D}$ for human subjects using the popular HRNet [64]  [^3] The configuration of the 2D keypoint estimation model can be found here.. Given the 2D keypoint annotations, we are also able to discard over part-occluded subjects. Specially, we only keep the subject with the average confidence of 2D keypoints over $0.45$. For segmentation, we feed the 2D bounding box and the 2D keypoints into the sam_vit_h sub-model to extract the foreground mask.
Stage III: Coarse SMPL and Camera Estimation. We first estimate SMPL and camera parameters frame by frame, using a state-of-the-art single-image-based human pose and shape estimation method PARE [36]. To achieve better performance in complex scenes in the wild, we adapt the model pre-trained on the in-the-wild 3DPW dataset [68]. The PARE infers human body pose/shape parameters ($\theta$/$\beta$), and the global camera parameters (rotation matrix $R$ and the 3D offset $T$).
To retain the remaining video clips with considerable viewpoint shifts and human movements, we discard the clips with viewpoint angle changes lower than $\frac{\pi}{4}~{}rad$. And we automatically select $N=20$ most non-trivial frames, which keep the pose diversity to the greatest extent possible.
As mentioned in the main paper, we double-check the consistency between the SAM masks and the SMPL masks. Intuitively, the SMPL mask denotes the naked body, while the SAM mask contains the clothed body. Therefore, the SMPL mask should be mostly covered by the SAM mask (See Fig. A1 (a)  ~  (d)). We discard the subjects whose SAM masks are over 3×3$\times$3 × larger than their SMPL masks (See Fig. A1 (e)  ~  (h)). We also discard the subjects whose $10\%$ SMPL mask pixels are not covered by the SAM mask (See Fig. A1 (i)  ~  (l)).
Stage IV: Refining SMPL and Camera In-the-loop. For high-quality annotations, here we refine the coarse SMPL parameters ($\theta,\beta$) and camera parameters ($R,T$) obtained in Stage I. To achieve temporally smooth results, we regularize the differences in parameters between adjacent frames, which is given by
$\mathcal{L}^{s}_{\theta}=\sum_{i=1}^{N-1}\left\|\theta^{i}-\theta^{i+1}\right\|_{2},\mathcal{L}^{s}_{R}=\sum_{i=1}^{N-1}\left\|R^{i}-R^{i+1}\right\|_{2},\mathcal{L}^{s}_{T}=\sum_{i=1}^{N-1}\left\|T^{i}-T^{i+1}\right\|_{2},$ (1)
$\mathcal{L}^{s}_{2D}=\sum_{i=1}^{N-1}\left\|\Pi(J_{3D}(\theta^{i},\beta^{i});R^{i},T^{i})-\Pi(J_{3D}(\theta^{i+1},\beta^{i+1});R^{i+1},T^{i+1})\right\|_{2},$
where $J_{3D}(\theta,\beta)$ infers the 3D keypoints of the human body, and the $\Pi$ denotes the 2D projection, and $i$ denotes the $i_{th}$ frame of the input video. Notice that the body shapes ($\beta$) are treated as constants across the input video.
Figure A1: SMPL and SAM consistency. The green/red borders denote good/bad outputs, respectively.
In addition, we align the human body parameters to the 2D keypoints and SAM masks, which are given by
$\mathcal{L}_{2D}=\sum_{i=1}^{N}\left\|\Pi(J_{3D}(\theta^{i},\beta^{i}))-J^{i}_{2D}\right\|_{2},\mathcal{L}_{\mathrm{mask}}=\sum_{i=1}^{N}d_{CD}(\Pi(V_{3D}(\theta^{i},\beta^{i})),M^{i}),$ (2)
where $d_{CD}$ denotes the chamfer distance between point clouds, and the $V_{3D}(\theta,\beta)$ infers the 3D vertex of the human body. The $M^{i}$ denotes the foreground d pixels cloud of the $i$ frame. We also regularize $\theta$ to avoid out-of-domain poses [7], using the Gaussian Mixture Model (GMM) prior [50]
$\mathcal{L}_{\mathrm{prior}}=\sum_{i=1}^{N}\left\|GMM(\theta^{i})\right\|_{2}.$ (3)
We adopt the loss functions as mentioned above for supervision:
$\mathcal{L}=\lambda^{s}_{\theta}\mathcal{L}^{s}_{\theta}+\lambda^{s}_{R}\mathcal{L}^{s}_{R}+\lambda^{s}_{T}\mathcal{L}^{s}_{T}+\lambda^{s}_{2D}\mathcal{L}^{s}_{2D}$ (4)
$+\lambda_{2D}\mathcal{L}_{2D}+\lambda_{\mathrm{mask}}\mathcal{L}_{\mathrm{mask}}+\lambda_{\mathrm{prior}}\mathcal{L}_{\mathrm{prior}}.$
We empirically set the loss weights as $\lambda^{s}_{\theta}=100$, $\lambda^{s}_{R}=1000$, $\lambda^{s}_{T}=50$, $\lambda^{s}_{2D}=100$, $\lambda_{2D}=100$, $\lambda_{\mathrm{mask}}=100$ and $\lambda_{\mathrm{prior}}=0.1$. We adopt the LBFGS [61] optimizer with the learning rate $lr=1.0$. We optimize for $20$ steps on a single Geforce GTX 1080Ti GPU.
Qualitative comparisons between coarse and refined SMPL and camera parameters can be found in Fig. A2.
Figure A2: Comparison of the coarse and refined SMPL parameters.
Appendix B Additional Links
We release the WildAvatar dataset with a permanent doi 10.5281/zenodo.11526806. We also provide the Croissant metadata. The authors bear all responsibility in case of violation of rights, and confirm that this dataset is open-sourced under the CC BY 4.0 International license.